Find the intervals where 1 f(x) = 3x1/3 x + 1 is concave down. Select one: 0 (-0,0) 0 (-00,00) (2/3,0) 0 (0,0)
Answer #1

1 Given f (x ) = 3x3 ---x+1 4 1 3(x) = ( 3*-*x+1) - 1 (1) +0 -1 = 3 با بيا = X 4 $(x) = - X -0 ما با II X F(x) is not defi